OpenStudy (anonymous):

Which of the numbers below are solutions for the equation (x + 7)(x - 8) = 0 ? Choose all of the correct answers below. -7 7 -8 8

OpenStudy (anonymous):

-7 and 8

OpenStudy (anonymous):

How you know ?

OpenStudy (anonymous):

(x+7)(x-8)=0 x+7=0 ; x=-7 x-8=0 ; x = 8

OpenStudy (anonymous):

How you know ?

OpenStudy (anonymous):

ok .